The bob of mass 15kg is moving with a constant velocity of 5m/s in a horizontal plane at the end of the cord 8 meters length. Determine the tension in the cord and the period ( time required to complete one revolution).
1
Expert's answer
2018-08-21T10:21:09-0400
ω=v/R=5/8=0.625 s^(-1)
T=2π/ω=(2·3.14)/0.625=10.048 s
T_thread=ma
a=v^2/R
T_thread=(m·v^2)/R=(15·5^2)/8=46.875 N